# The Gauss map of a projective foliation

### v2

@@ Abstract.
The $1^{st}$ author was partially supported by Universidad de Guanajuato and Université de Rennes 1.
@@ 1. Introduction
In a certain sense, this article gives us a generalization in the context of foliation theory of a classic concept in algebraic geometry, we refer to projective duality. The study of projective duality was started by Monge and his student Poncelet and later developed by Plücker, who was interested in the study of the dual curve of a plane curve (that is, the curve defined by the tangent lines). He proved that the degree of the dual curve $\breve{\Gamma}$ of a curve $\Gamma$ satisfies: $\deg{\breve{\Gamma}}=\deg{\Gamma}(\deg{\Gamma}-1)$, when the curve is smooth. For the singular case, the nature of the singularities of the curve must be taken into account: in this case, we have the famous Plücker’s formulas. He also studied self-dual curves, which are those that are projectively equivalent to their dual. In the case of hypersurfaces with isolated singularities in a projective space $\mathbb{P}^{n}$, it is necessary to understand the works by Katz and Laumon to obtain significant results. For the smooth case, we have an analogous formula to the previous one: $\deg{\breve{H}}=\deg{H}(\deg{H}-1)^{n-1}$, result that Laumon attributes to Katz. The case of hypersurfaces with singularities is more subtle, for which we refer to the work of Laumon (see [22]). In this work he uses the invariants developed by Teissier in [33]. These invariants generalize the classical Milnor number of singularities.
@@
As we will see later, we can associate to a foliation $\mathcal{F}$ of codimension 1 of the complex projective space $\mathbb{P}^{n}$, its Gauss map. This is a rational map that sends to a non-singular point $m$, the hyperplane tangent to the leaf that passes through $m$. The space of codimension 1 foliations in $\mathbb{P}^{n}$ for $n>2$ is not completely known. Only spaces of degrees 00, $1$ and $2$ in $\mathbb{P}^{n}$, for $n\geq 3$ have been fully described (see [9]). The case of degree 3 seems approachable after works [10] and [13].
@@
The main objective of this work is to give some results on the Gauss map of codimension one foliations in $\mathbb{P}^{n}$, with special attention to the case where the map is birational. As well as a large number of examples where we make explicit calculations to clarify what happens in each case. This topic has been studied by some authors such as Fassarella and Pereira in [17], [18], and [19].
@@
The article is structured as follows, in section 2 we give some statements of structure, and we give a list of open problems that we find interesting. In section 3, we can find the proofs of the results. Finally, in section 4 we study the exceptional component of the space of foliations of codimension 1 in $\mathbb{P}^{n}$.
@@
Acknowledgement: We would like to thank Serge Cantat and Julie Deserti for helpful discussions. We would also like to thank Thiago Fassarella for his valuable comments and suggestions on this article.
@@ 2. Basic properties, Results, and Examples
In order not to write many subscripts, we will adopt the following convention for the notation of the coordinates: If we are working in a projective space $\mathbb{P}^{n}$ with $n=2$, $n=3$ and $n=4$ we will use $(x,y,z,t,s)$. For general n we will use $(z_{0},...,z_{n})$. Except in section 3.7 where, due to the nature of the proof of theorem 4, it is more useful to work with $(z_{1},z_{2},z_{3},z_{4})$.
Let $\mathcal{F}$ be a codimension one holomorphic foliation on a $n$-dimensional complex manifold $X$, $n\geq 2$, with singular set $Sing(\mathcal{F})$. This foliation defines a codimension one distribution on $X\setminus Sing(\mathcal{F})$ denoted as $\mathcal{D}_{\mathcal{F}}$: given $z\notin Sing(\mathcal{F})$ then $\mathcal{D}_{\mathcal{F}}(z)=$ the tangent space at $z$ of the leaf of $\mathcal{F}$ through $z$. We will always assume that $cod(Sing(\mathcal{F}))\geq 2$.
@@
In other words, if we fix an affine coordinate system $z=(z_{1},...,z_{n})\in\mathbb{C}^{n}\subset\mathbb{P}^{n}$ then $\mathcal{F}$ is defined by some polynomial integrable 1-form, say $\omega=\sum_{j=1}^{n}P_{j}(z)dz_{j}$. In this affine system, we have
@@
In particular, if $z_{o}=(z_{o1},...,z_{on})\in\mathbb{C}^{n}\setminus Sing(\mathcal{F})$ then $\mathcal{G}_{\mathcal{F}}(z_{o})=$ the plane which in this affine system is given by $\ell(z)=0$, where
@@
The map $\mathcal{G}_{\mathcal{F}}$ can be expressed also in homogeneous coordinates. This is done as follows: let $\Pi\colon\mathbb{C}^{n+1}\setminus\{0\}\to\mathbb{P}^{n}$ be the canonical projection. It is known that the foliation $\Pi^{*}(\mathcal{F})$ on $\mathbb{C}^{n+1}\setminus\{0\}$ can be extended to a foliation on $\mathbb{C}^{n+1}$, say $\widetilde{\mathcal{F}}$. This foliation can be defined by a polynomial integrable 1-form, say $\widetilde{\omega}=\sum_{j=0}^{n}Q_{j}(z)dz_{j}$, where the coefficients $Q_{j^{\prime}s}$ are all homogeneous polynomials of the same degree $d=degre(\mathcal{F})+1$ and such that $i_{R}\,\widetilde{\omega}=0$, where $R=\sum_{j=0}^{n}z_{j}\frac{\partial}{\partial z_{j}}$, is the radial vector field on $\mathbb{C}^{n+1}$ and $i_{R}$ denotes the interior product by $R$.
@@
Note also that $\Pi(Sing(\widetilde{\mathcal{F}}))$ is the indetermination set of $\mathcal{G}_{\mathcal{F}}$.
We say that $\mathcal{G}_{\mathcal{F}}$ is dominant if it is generically of maximal rank.
Remark 2.1.
If $\mathcal{G}_{\mathcal{F}}=[Q_{0}:...:Q_{n}]$ in homogeneous coordinates, then $\mathcal{G}_{\mathcal{F}}$ is dominant if, and only if, $dQ_{0}\wedge dQ_{1}\wedge...\wedge dQ_{n}\not\equiv 0$.
@@
Let $R$ be the radial vector field on $\mathbb{C}^{n+1}$. Since $Q_{0},...,Q_{n}$ are homogeneous of the same degree, say $k$, by Euler’s identity we have
$i_{R}[dQ_{0}\wedge dQ_{1}\wedge...\wedge dQ_{n}]=k\sum_{j=0}^{n}(-1)^{j}Q_{j}\,dQ_{0}\wedge...\wedge\widehat{dQ_{j}}\wedge...\wedge dQ_{n}:=\Theta$
In the above formula the symbol $\widehat{dQ_{j}}$ means the omission of $dQ_{j}$ in the product. This implies that
$\Theta=i_{R}[J(z)dz_{0}\wedge...\wedge dz_{n}]=J(z)\sum_{j=0}^{n}(-1)^{j}z_{j}\,dz_{0}\wedge...\wedge\widehat{dz_{j}}\wedge...\wedge dz_{n}:=J(z)\Gamma$
On the other hand, if $z\in Sing(\mathcal{F})$ then $\Theta(z)=0$ which implies $J(z)=0$, because $\Gamma(z)\neq 0$.∎
@@
The proof is based in the fact that $m\in\mathcal{U}$ if, and only if, the derivative, $D\mathcal{G}_{\mathcal{F}}(m)\colon T_{m}\mathbb{P}^{n}\to T_{\mathcal{G}_{\mathcal{F}}(m)}\breve{\mathbb{P}}^{n}$ is an isomorphism. This computation is straightforward and is left to the reader.
As a consequence, if $\mathcal{F}=L^{*}(\mathcal{F}^{\prime})$ where, $\mathcal{F}^{\prime}$ is a codimension one foliation on $\mathbb{P}^{m}$ with $m<n$, and $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{m}$ is a linear projection, then the Gauss map $\mathcal{G}_{\mathcal{F}}$ cannot be dominant. In fact, if $z\in\mathbb{P}^{n}\setminus Sing(\mathcal{F})$ then the leaf $\mathcal{L}_{z}$ of $\mathcal{F}$ through $z$ contains a straight line through $z$ such that $\mathcal{F}$ is tangent to $\mathcal{L}_{z}$ along the line. In particular, the tangency with the tangent hyperplane cannot be of Morse type.
@@
Remark 2.3.
We would like to observe that the map $F\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{1}$ of (1) in the statement of proposition 1 is birationally equivalent to the map $G(x^{\prime},y^{\prime},z^{\prime},t^{\prime})=z^{\prime}/t^{\prime}$. In particular, the closure of the levels of $F$ in $\mathbb{P}^{3}$ are rational surfaces.
@@
Let $\mathbb{F}ol(d;n)$ be the space of foliations of degree $d$ on $\mathbb{P}^{n}$. Given $\mathcal{F}\in\mathbb{F}ol(d;n)$ denote its Gauss map as $\mathcal{G}_{\mathcal{F}}$. The topological degree of $\mathcal{G}_{\mathcal{F}}$ is defined as $tdg(\mathcal{G}_{\mathcal{F}})=\#(\mathcal{G}_{\mathcal{F}}^{-1}(m))$, where $m$ is a regular value of $\mathcal{G}_{\mathcal{F}}$ (see [30]). For instance, if $\mathcal{G}_{\mathcal{F}}$ is not dominant then $tdg(\mathcal{G}_{\mathcal{F}})=0$, because a regular value of $\mathcal{G}_{\mathcal{F}}$ is not in the image.
If $\mathcal{G}_{\mathcal{F}}$ is dominant then $tdg(\mathcal{G}_{\mathcal{F}})<+\infty$. Set
@@
and $\mathcal{M}(d,n)=max(\mathcal{N}(d,n))$. In proposition 2.6 of Fassarella’s thesis (see [19]) we can find a close version of the following result. For the sake of completeness, we include the proof of this in §3.1.
@@
Remark 2.5.
A natural problem is to determine exactly the set $\mathcal{N}(d,n)$. For instance, $\mathcal{N}(2,3)=\{0,1,2,3,4\}$. This is a consequence of the classification of degree two foliations of $\mathbb{P}^{3}$ (see [9] and example 9). In example 11 we will see that $\mathcal{N}(3,3)=\{0,1,2,3,4,5,6,7,8,9\}$ and in remark 2.7 that $1\in\mathcal{N}(d,3)$, for all $d\geq 1$.
@@
Remark 2.6.
If $\mathcal{F}$ is a foliation in $\mathbb{P}^{n}$ with more than one Morse singularity then $\mathcal{G}_{\mathcal{F}}$ cannot be birational. We leave the details of the proof for the reader.
For instance, if $\mathcal{F}$ has the first integral, given in affine coordinate by $f(x,y,z)=x^{2}+y^{2}+z^{2}-x^{3}$, then it has two Morse singularities in $\mathbb{C}^{3}$; the points $(0,0,0)$ and $(2/3,0,0)$, so that $\mathcal{G}_{\mathcal{F}}$ is not birational.
@@
If $n\geq 2$ then a birational map $\mathcal{G}\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{n}$, which is not an automorphism, always contracts a hypersurface of $\mathbb{P}^{n}$; there exists a hypersurface $V\subset\mathbb{P}^{n}$ such that $\mathcal{G}(V)$ is an algebraic subset of $\mathbb{P}^{n}$ of dimension $<n-1$.
@@
Conversely, if the hypersurface $(h=0)$ is contracted by $\mathcal{G}$ and $h$ is irreducible then $h\,|\,f$.
@@
Corollary 2.1.
If the Gauss map $\mathcal{G}_{\mathcal{F}}$ of a foliation $\mathcal{F}$ on $\mathbb{P}^{n}$ is birational then it always contracts some hypersurface $V\subset\mathbb{P}^{n}$. Moreover, $V$ is contracted to a point of ˘$\breve{\mathbb{P}}^{n}$ if, and only if $V$ is an invariant hyperplane.
@@
Let $\mathcal{F}$ be a foliation on $\mathbb{P}^{n}$ with birational Gauss map $\mathcal{G}$. Then the Gauss map of the transformed foliation $\mathcal{G}_{*}(\mathcal{F})$ is also birational.
@@
On the other hand, if $\mathcal{F}$ is a foliation on $\mathbb{P}^{2}$ of degree $d\geq 1$ then the Gauss map $\mathcal{G}_{\mathcal{F}}$ is generically $d$ to one. Therefore $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{2}\dasharrow\breve{\mathbb{P}}^{2}$ is birational if, and only if $d=1$.
In homogeneous coordinates, a degree one foliation in $\mathbb{P}^{2}$ is defined by an action of $\mathbb{C}^{2}$ on $\mathbb{C}^{3}$ generated by the radial vector field $R=\sum_{j=0}^{2}z_{j}\frac{\partial}{\partial z_{j}}$ and a linear vector field $L$ of $\mathbb{C}^{3}$ such that $R\wedge L\neq 0$. In the generic case, after a linear change of variables, we can suppose that $L=\sum_{j=0}^{2}\lambda_{j}z_{j}\frac{\partial}{\partial z_{j}}$, where $\lambda_{i}\neq\lambda_{j}$ if $i\neq j$.
@@
The case of foliations of degree one on $\mathbb{P}^{n}$, $n\geq 3$. It is known (see [21]) that the space of foliations of degree one in $\mathbb{P}^{n}$, $n\geq 3$, has two irreducible components:
1. The logarithmic component $\mathcal{L}(1,1,1;n)$ whose generic element is, as we said before, defined in homogeneous coordinates by a 1-form linearly equivalent to the following:
@@
2. The component $\mathcal{L}(2,1;n)$, whose elements have a rational first integral of the form $\frac{q}{\ell^{2}}$, where $q$ is irreducible of degree two and $\ell$ of degree one.
In the generic case, the Gauss map of the associated foliation is birational and not a linear isomorphism from $\mathbb{P}^{n}$ to ˘$\breve{\mathbb{P}}^{n}$. In fact, in this case, we can assume that in some homogeneous coordinates $z=(z_{0},z_{1},...,z_{n})$ we have $q(z)=\sum_{j=1}^{n}z_{j}^{2}$ and $\ell(z)=z_{0}$. The foliation is defined by the form

### v3

@@ Abstract.
The $1^{st}$ author was partially supported by Universidad de Guanajuato, Université de Rennes 1 and LaSol International Research Laboratory.
@@ 1. Introduction
In a certain sense, this article gives us a generalization in the context of foliation theory of a classic concept in algebraic geometry, we refer to projective duality. The study of projective duality was started by Monge and his student Poncelet and later developed by Plücker, who was interested in the study of the dual curve of a plane curve (that is, the curve defined by the tangent lines). He proved that the degree of the dual curve $\breve{\Gamma}$ of a curve $\Gamma$ satisfies: $\deg{\breve{\Gamma}}=\deg{\Gamma}(\deg{\Gamma}-1)$, when the curve $\Gamma$ is smooth. For the singular case, the nature of the singularities of the curve must be taken into account: in this case, we have the famous Plücker’s formulas. He also studied self-dual curves, which are those that are projectively equivalent to their dual. In the case of hypersurfaces with isolated singularities in a projective space $\mathbb{P}^{n}$, it is necessary to understand the works by Katz and Laumon to obtain significant results. For the smooth case, we have an analogous formula to the previous one: $\deg{\breve{H}}=\deg{H}(\deg{H}-1)^{n-1}$, result that Laumon attributes to Katz. The case of hypersurfaces with singularities is more subtle, for which we refer to the work of Laumon (see [24]). In this work he uses the invariants developed by Teissier in [34]. These invariants generalize the classical Milnor number of singularities. While these topics still offer many directions, our main intention in this work is to obtain analogous results in the context of codimension one foliations.
@@
As we shall see later, we can associate to a foliation $\mathcal{F}$ of codimension 1 of the complex projective space $\mathbb{P}^{n}$, its Gaussian map. This is a rational map that sends to a non-singular point $m$, the hyperplane tangent to the leaf that passes through $m$. The space of codimension 1 foliations in $\mathbb{P}^{n}$ for $n>2$, is not completely known. Only spaces of degrees 0, 11 and $2$ in $\mathbb{P}^{n}$, for $n\geq 3$ have been fully described (see [10]). The case of degree 3 seems approachable after the works [11] and [15]. Given the limited understanding of these spaces, the Gauss map of a foliation could be a useful tool for the study of these spaces.
@@
Motivated by this, in the present work we carry out a detailed study of this map. Concretely, the main objective of this article is to give some results on the Gauss map of codimension one foliations in $\mathbb{P}^{n}$, with special attention to the case where the map is birational. In addition, there are a large number of examples in which we make explicit calculations to clarify what happens in each case. This topic has been studied by some authors, such as Fassarella and Pereira in [19], [20], and [21].
With the aim of offering a more insightful first reading, we describe in more detail the main results of the article.
We begin with Theorem 1, which asserts that the maximum topological degree (i.e., the cardinality of a generic fiber of the Gauss map) that a foliation of $\mathbb{P}^{3}$ of degree $d$ can attain is $d^{2}$, and when this occurs, the foliation admits a rational first integral of the form $\frac{F}{L^{d+1}}$, where $F$ has degree $d+1$ and $L$ has degree $1$. In Theorem 2, we prove that if $F\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{2}$ is a rational map whose fibers are not plane curves, except for a finite number of them and there is a foliation $\mathcal{G}$ of $\mathbb{P}^{2}$ with the property that the foliation $\mathcal{F}=F^{*}(\mathcal{G})$ has a birational Gauss map and a contracted hypersurface that is not invariant, then $\mathcal{F}$ has a rational first integral. Theorem 3 provides, among other things, a characterization of foliations in $\mathbb{P}^{3}$ for which the Gauss map is birational. This characterization depends on the geometry of their singular set. Finally, Theorem 4 deals with monomial foliations in $\mathbb{P}^{3}$ and shows that if the Gauss map is dominant, then the foliations are of either a special logarithmic type or a specific pull-back from $\mathbb{P}^{2}$. In both cases, the Gauss map is birational.
Based on these results, we present several corollaries and numerous examples that illustrate, in some cases, that the hypotheses are sufficient but not necessary. Additionally, we study in detail the Gauss map of the foliations in the exceptional component constructed in [10]. We also describe the algorithm we have used to construct the inverse of a Gauss map, whenever it exists.
@@
The article is structured as follows, in Section 2 we give the statements of structure, and we give a list of open problems that we find interesting. In Section 3, we can find the proofs of the results. Finally, in Section 4 we study the exceptional component of the space of foliations of codimension 1 in $\mathbb{P}^{n}$.
@@
Acknowledgement: We would like to thank Serge Cantat and Julie Deserti for helpful discussions. We also thank Thiago Fassarella and Jorge V. Pereira for valuable comments and suggestions on this work. The first author is very grateful for the hospitality received at the University of Rennes. Finally, the authors are grateful to anonymous referees whose comments helped improve this article.
@@ 2. Basic properties, Results, and Examples
In order not to write many subscripts, we will adopt the following convention for the notation of the coordinates: If we are working in a projective space $\mathbb{P}^{n}$ with $n=2$, $n=3$ and $n=4$ we will use $(x,y,z,t,s)$. For general $n$, we will use $(z_{0},...,z_{n})$. Except in Section 3.7 where, due to the nature of the proof of Theorem 4, it is more useful to work with $(z_{1},z_{2},z_{3},z_{4})$.
Let $\mathcal{F}$ be a codimension one holomorphic foliation on a $n$-dimensional complex manifold $X$, $n\geq 2$, with singular set $Sing(\mathcal{F})$. This foliation defines a codimension one distribution on $X\setminus Sing(\mathcal{F})$, denoted by $\mathcal{D}_{\mathcal{F}}$: given $z\notin Sing(\mathcal{F})$ then $\mathcal{D}_{\mathcal{F}}(z)$ is the tangent space at $z$ of the leaf of $\mathcal{F}$ through $z$. We will always assume that $cod(Sing(\mathcal{F}))\geq 2$.
@@
In other words, if we fix an affine coordinate system $z=(z_{1},...,z_{n})\in\mathbb{C}^{n}\subset\mathbb{P}^{n}$ then $\mathcal{F}$ is defined by some polynomial integrable 1-form, say $\omega=\sum_{j=1}^{n}P_{j}(z)dz_{j}$. In this affine system, we have the following.
@@
In particular, if $z_{o}=(z_{o1},...,z_{on})\in\mathbb{C}^{n}\setminus Sing(\mathcal{F})$ then $\mathcal{G}_{\mathcal{F}}(z_{o})$ is the plane which, in this affine system, is given by $\ell(z)=0$, where
@@
The map $\mathcal{G}_{\mathcal{F}}$ can be expressed also in homogeneous coordinates. This is done as follows: Let $\Pi\colon\mathbb{C}^{n+1}\setminus\{0\}\to\mathbb{P}^{n}$ be the canonical projection. It is known that the foliation $\Pi^{*}(\mathcal{F})$ on $\mathbb{C}^{n+1}\setminus\{0\}$ can be extended to a foliation on $\mathbb{C}^{n+1}$, say $\widetilde{\mathcal{F}}$. This foliation can be defined by an integrable polynomial 1-form, say $\widetilde{\omega}=\sum_{j=0}^{n}Q_{j}(z)dz_{j}$, where the coefficients $Q_{j^{\prime}s}$ are all homogeneous polynomials of the same degree $d+1$, which is the degree of the foliation plus one, and such that $i_{R}\,\widetilde{\omega}=0$, where $R=\sum_{j=0}^{n}z_{j}\frac{\partial}{\partial z_{j}}$, is the vector radial field in $\mathbb{C}^{n+1}$ and $i_{R}$ denotes the interior product by $R$.
@@
Note also that $\Pi(Sing(\widetilde{\mathcal{F}})\setminus\{0\})$ is the indeterminacy set of $\mathcal{G}_{\mathcal{F}}$. Now we summarize the above in the following definition, which we will use throughout the article.
Definition 1.
Let $\mathcal{F}$ be a codimension one holomorphic foliation on $\mathbb{P}^{n}$, given by the 1-form $\omega=\sum_{j=0}^{n}Q_{j}(z)dz_{j}$. The Gauss map of $\mathcal{F}$ is defined as the rational function:
$\mathcal{G}_{\mathcal{F}}:\mathbb{P}^{n}$ $\dashrightarrow\breve{\mathbb{P}}^{n}$
$z$ $\mapsto[Q_{0}(z):...:Q_{n}(z)].$
As usual, we will say that $\mathcal{G}_{\mathcal{F}}$ is dominant if it is generically of maximal rank.
Remark 1.
If $\mathcal{G}_{\mathcal{F}}=[Q_{0}:...:Q_{n}]$ in homogeneous coordinates, then $\mathcal{G}_{\mathcal{F}}$ is dominant if and only if $dQ_{0}\wedge dQ_{1}\wedge...\wedge dQ_{n}\not\equiv 0$.
@@
Let $R$ be the radial vector field on $\mathbb{C}^{n+1}$. Since $Q_{0},...,Q_{n}$ are homogeneous of the same degree, say $d$, by Euler’s identity we have
$i_{R}[dQ_{0}\wedge dQ_{1}\wedge...\wedge dQ_{n}]=d\sum_{j=0}^{n}(-1)^{j}Q_{j}\,dQ_{0}\wedge...\wedge\widehat{dQ_{j}}\wedge...\wedge dQ_{n}:=\Theta.$
In the above formula the symbol $\widehat{dQ_{j}}$ means the omission of $dQ_{j}$ in the product.
This implies that
$\Theta=i_{R}[J(z)dz_{0}\wedge...\wedge dz_{n}]=J(z)\sum_{j=0}^{n}(-1)^{j}z_{j}\,dz_{0}\wedge...\wedge\widehat{dz_{j}}\wedge...\wedge dz_{n}:=J(z)\Gamma.$
On the other hand, if $z\in Sing(\mathcal{F})$ then $\Theta(z)=0$, which implies $J(z)=0$, because $\Gamma(z)\neq 0$.∎
@@
The proof is based on the fact that $m\in\mathcal{U}$ if, and only if, the derivative $D\mathcal{G}_{\mathcal{F}}(m)\colon T_{m}\mathbb{P}^{n}\to T_{\mathcal{G}_{\mathcal{F}}(m)}\breve{\mathbb{P}}^{n}$ is an isomorphism. In fact, at a Morse tangency, the foliation is given by a form with linear part equivalent to:
$dx_{n}+x_{1}dx_{1}+...+x_{n-1}dx_{n-1},$
and it is immediate to verify that $D\mathcal{G}_{\mathcal{F}}(m)$ is an isomorphism. Conversely, if $m$ is not a Morse tangency, the form defining the foliation at $m$ has linear part at $m$ equivalent to: $dx_{n}+dq$, where $q$ is a quadratic form of rank less than n1n-1, so that $D\mathcal{G}_{\mathcal{F}}(m)$ cannot be an isomorphism. Finally, this implies that $D\mathcal{G}_{\mathcal{F}}(\mathcal{U})$ is Zariski open and dense if, and only if, $\mathcal{U}$ is also.
As a consequence, if $\mathcal{F}=L^{*}(\mathcal{F}^{\prime})$, where $\mathcal{F}^{\prime}$ is a codimension one foliation on $\mathbb{P}^{m}$ with $m<n$, and $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{m}$ is a linear projection, then the Gauss map $\mathcal{G}_{\mathcal{F}}$ cannot be dominant. In fact, if $z\in\mathbb{P}^{n}\setminus Sing(\mathcal{F})$ then the leaf $\mathcal{L}_{z}$ of $\mathcal{F}$ through $z$ contains a straight line through $z$ such that $\mathcal{F}$ is tangent to $\mathcal{L}_{z}$ along the line. In particular, the tangency with the tangent hyperplane cannot be of Morse type.
@@
Remark 3.
We would like to observe that the map $F\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{1}$ of (1), in the statement of proposition 1 is birationally equivalent to the map $G(x^{\prime},y^{\prime},z^{\prime},t^{\prime})=z^{\prime}/t^{\prime}$. In particular, the closure of the levels of $F$ in $\mathbb{P}^{3}$ are rational surfaces.
@@
Let $\mathbb{F}ol(d;n)$ be the space of foliations of codimension 1 of degree $d$ on $\mathbb{P}^{n}$ (we use the definition given in section 1 of [10]). Given $\mathcal{F}\in\mathbb{F}ol(d;n)$ denote its Gauss map as $\mathcal{G}_{\mathcal{F}}$. The topological degree of $\mathcal{G}_{\mathcal{F}}$ is defined as $tdg(\mathcal{G}_{\mathcal{F}})=\#(\mathcal{G}_{\mathcal{F}}^{-1}(m))$, where $m$ is a regular value of $\mathcal{G}_{\mathcal{F}}$ (see [31]). For instance, if $\mathcal{G}_{\mathcal{F}}$ is not dominant then $tdg(\mathcal{G}_{\mathcal{F}})=0$, because a regular value of $\mathcal{G}_{\mathcal{F}}$ is not in the image.
On the other hand, if $\mathcal{G}_{\mathcal{F}}$ is dominant then $tdg(\mathcal{G}_{\mathcal{F}})<+\infty$, therefore we can define
@@
and $\mathcal{M}(d,n)=max(\mathcal{N}(d,n))$. In Section 2.1 of Fassarella’s thesis (see [20]) several results on the topological degree are established; for example, it is shown that the map assigning to a foliation its topological degree is upper semicontinuous (Proposition 2.1). Proposition 2.6, on the other hand, contains a statement close to the one presented below. For the sake of completeness, we include the complet proof of this in §3.1.
@@
Remark 5.
A natural problem is to determine exactly the set $\mathcal{N}(d,n)$. For instance, $\mathcal{N}(2,3)=\{0,1,2,3,4\}$. This is a consequence of the classification of degree two foliations of $\mathbb{P}^{3}$ (see [10] and example 10). In example 11 we will see that $\mathcal{N}(3,3)=\{0,1,2,3,4,5,6,7,8,9\}$ and in remark 7 that $1\in\mathcal{N}(d,3)$, for all $d\geq 1$.
@@
Another interesting example is provided by the Fermat foliation of degree $d$ on $\mathbb{P}^{n}$, given by the 1-form:
$\omega=\big(\sum_{i=1}^{n}z_{i}^{d+1}\big)dz_{0}-z_{0}z_{1}^{d}dz_{1}-...-z_{0}z_{n}^{d}dz_{n},$
where it can be seen that the topological degree of the Gauss map is $d^{n-1}$.
Let $m$ be a singular point of a foliation $\mathcal{F}\in\mathbb{F}ol(d;n)$. We will say that $m$ is a Morse singularity for the foliation, if it admits a holomorphic first integral of the form $f=Q+h.o.t.$, where $Q$ is a non-degenerate quadratic form.
Remark 6.
If $\mathcal{F}$ is a foliation in $\mathbb{P}^{n}$ with more than one Morse singularity then $\mathcal{G}_{\mathcal{F}}$ cannot be birational.
For the proof, take $P$ a linear hyperplane through $m$ transversal to $\mathcal{F}$ (i.e., the restriction of $Q$ to $P$ is non-degenerate), then the restriction of $\mathcal{F}$ to $P$ has a Morse singularity at $m$. Furthermore, if $P^{\prime}$ is a hyperplane sufficiently close to $P$, the restriction of $\mathcal{F}$ to $P^{\prime}$ has a Morse singularity $m^{\prime}$ near $m$.
Now suppose that $\mathcal{F}$ has two Morse singularities $a$ and $b$. Let $P$ be a hyperplane containing both, transversal to $\mathcal{F}$ at these points. If $P^{\prime}$ is a nearby hyperplane not containing $a$ and $b$, then $\mathcal{F}$ has no singularities near $a$ and $b$, since Morse singularities are isolated. However, the restriction $\mathcal{F}_{|P^{\prime}}$ has two distinct Morse singularities. It follows that the Gauss map of $\mathcal{F}$ takes the same value $P^{\prime}$ at both points. Consequently, the Gauss map is not birational.
For instance, if $\mathcal{F}$ has the first integral, given in affine coordinate by $f(x,y,z)=x^{2}+y^{2}+z^{2}-x^{3}$, then it has two Morse singularities in $\mathbb{C}^{3}$, namely, the points $(0,0,0)$ and $(2/3,0,0)$, so that $\mathcal{G}_{\mathcal{F}}$ is not birational.
@@
If $n\geq 2$ then a birational map $\mathcal{G}\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{n}$, which is not an automorphism, always contracts a hypersurface of $\mathbb{P}^{n}$. That is, there exists a hypersurface $V\subset\mathbb{P}^{n}$ such that $\mathcal{G}(V)$ is an algebraic subset of $\mathbb{P}^{n}$ of dimension <n1<n-1.
@@
Conversely, if the hypersurface $(h=0)$ is contracted by $\mathcal{G}$ and $h$ is irreducible then h|fh\,|\,f.
@@
Corollary 1.
If the Gauss map $\mathcal{G}_{\mathcal{F}}$ of a foliation $\mathcal{F}$, of codimension one on $\mathbb{P}^{n}$ is birational then it always contracts some hypersurface $V\subset\mathbb{P}^{n}$. Moreover, $V$ is contracted to a point of $\breve{\mathbb{P}}^{n}$ if, and only if, $V$ is an invariant hyperplane.
@@
Let $\mathcal{F}$ be a foliation, of codimension one on $\mathbb{P}^{n}$ with birational Gauss map $\mathcal{G}$. Then the Gauss map of the transformed foliation $\mathcal{G}_{*}(\mathcal{F})$ is also birational. Moreover, $\mathcal{G}_{\mathcal{G}_{*}(\mathcal{F})}=\mathcal{G}^{-1}$.
@@
On the other hand, if $\mathcal{F}$ is a foliation on $\mathbb{P}^{2}$ of degree $d\geq 1$ then the Gauss map $\mathcal{G}_{\mathcal{F}}$ is generically $d$ to one. Therefore, $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{2}\dasharrow\breve{\mathbb{P}}^{2}$ is birational if, and only if $d=1$. Indeed $\mathcal{N}(d,2):=\{d\}$.
In homogeneous coordinates, a degree one foliation on $\mathbb{P}^{2}$ is defined by an action of $\mathbb{C}^{2}$ on $\mathbb{C}^{3}$ generated by the radial vector field $R=\sum_{j=0}^{2}z_{j}\frac{\partial}{\partial z_{j}}$ and a linear vector field $L$ of $\mathbb{C}^{3}$, such that $R\wedge L\neq 0$. In the generic case, after a linear change of variables, we can suppose that $L=\sum_{j=0}^{2}\lambda_{j}z_{j}\frac{\partial}{\partial z_{j}}$, where $\lambda_{i}\neq\lambda_{j}$ if $i\neq j$.
@@
The case of foliations of degree one on $\mathbb{P}^{n}$, $n\geq 3$. It is known (see chapter 2 of [23]) that the space of foliations of degree one in $\mathbb{P}^{n}$, $n\geq 3$, has two irreducible components:
1. The logarithmic component $\mathcal{L}(1,1,1;n)$ whose generic elements are, as we said before, defined in homogeneous coordinates by 1-forms linearly equivalent to the following:
@@
2. The component $\mathcal{L}(2,1;n)$, whose elements have a rational first integral of the form $\frac{q}{\ell^{2}}$, where $q$ is irreducible of degree two and $\ell$ has degree one.
In the generic case, the Gauss map of the associated foliation is birational and not a linear isomorphism from $\mathbb{P}^{n}$ to $\breve{\mathbb{P}}^{n}$. In fact, in this case, we can assume that in some homogeneous coordinates $z=(z_{0},z_{1},...,z_{n})$ we have $q(z)=\sum_{j=1}^{n}z_{j}^{2}$ and $\ell(z)=z_{0}$. The foliation is defined by the form